OpenStudy (anonymous):

If f(x) = x^5 + x^3 +x find f-1 (3) how do I find the inverse function? I know its switching the x and y but how do I isolate y after?

OpenStudy (anonymous):

The inverse function is a function such that if f(a)=b then f-1(b)=a.

OpenStudy (anonymous):

So solve x^5+x^3+x=3, and you'll have your answer. Here it is solved: x^5+x^3+x-3=0 (x-1)(x^4+x^3+2x^2+2x+3)-0 x=1

OpenStudy (anonymous):

thus f-1(3)=1

OpenStudy (anonymous):

is the (x-1) a factor?